# Einstein connection of nonsymmetric pseudo-Riemannian manifold

### v1

@@ 6 The terms $\delta_{i}(X,Y,Z)$
Here, we explicitly present the terms $\delta_{i}(X,Y,Z)\ (1\leq i\leq 5)$, see Section 5). Their expressions were verified with the help of MAPLE program.
@@ Conclusion
We presented explicitly the Einstein connection for nonsymmetric pseudo-Riemannian, in particular, for weak almost Hermitian manifolds, satisfying the $f^{2}$-torsion condition. We also obtained the corresponding formulas for almost contact metric manifolds, assuming the $f^{2}$-torsion condition. We expressed the torsion in terms of $\nabla^{g}F$ and $dF$, and the additional $\widetilde{Q}-$terms, explained precisely how the weak almost Hermitian case differs from the almost Hermitian case. The presented identities provide a new tool for constructing examples and for studying classes of nonsymmetric pseudo-Riemannian manifolds, in particular, weak almost Hermitian manifolds with the usual Gray-Hervella classification, in Einstein’s nonsymmetric gravitational theory. In further research we will study weak (para-) metric $f$-manifolds, in particular, and weak (para-) contact metric manifolds, with an Einstein connection $\nabla$ satisfying the $f^{2}$-torsion condition for vectors orthogonal to the distribution $\ker f$.

### v2

@@ 5 The terms $\delta_{i}(X,Y,Z)$
Here, we present the terms $\delta_{i}(X,Y,Z)\ (1\leq i\leq 5)$, see Section 4. Their expressions were verified with the help of MAPLE program.
@@ Conclusion
We presented explicitly the Einstein connection for nonsymmetric pseudo-Riemannian, in particular, for weak almost Hermitian manifolds, satisfying the $f^{2}$-torsion condition. We also obtained the corresponding formulas for almost contact metric manifolds, assuming the $f^{2}$-torsion condition. We expressed the torsion in terms of $\nabla^{g}F$ and $dF$, and the additional $\widetilde{Q}-$terms, explained precisely how the weak almost Hermitian case differs from the almost Hermitian case. The presented identities provide a new tool for constructing examples and for studying classes of nonsymmetric pseudo-Riemannian manifolds, in particular, weak almost Hermitian manifolds with the usual Gray-Hervella classification, in Einstein’s nonsymmetric gravitational theory. In further research we will study weak (para-) metric $f$-manifolds, in particular, weak (para-) contact metric manifolds, with an Einstein connections or connections with totally skew-symmetric torsion, satisfying the $f^{2}$-torsion condition for vectors orthogonal to the distribution $\ker f$.